# Multi-Modal Multi-Task (M3T) Federated Foundation Models for Embodied AI: Potentials and Challenges for Edge Integration

### v1

@@ Abstract
As embodied AI systems become increasingly multi-modal, personalized, and interactive, they must learn effectively from diverse sensory inputs, adapt continually to user preferences, and operate safely under resource and privacy constraints. These challenges expose a pressing need for machine learning models capable of swift, context-aware adaptation while balancing model generalization and personalization. Here, two methods emerge as suitable candidates, each offering parts of these capabilities: Foundation Models (FMs) provide a pathway toward generalization across tasks and modalities, whereas Federated Learning (FL) offers the infrastructure for distributed, privacy-preserving model updates and user-level model personalization. However, when used in isolation, each of these approaches falls short of meeting the complex and diverse capability requirements of real-world embodied environments. In this vision paper, we introduce Federated Foundation Models (FFMs) for embodied AI, a new paradigm that unifies the strengths of multi-modal multi-task (M3T) FMs with the privacy-preserving distributed nature of FL, enabling intelligent systems at the wireless edge. We collect critical deployment dimensions of FFMs in embodied AI ecosystems under a unified framework, which we name “EMBODY”: Embodiment heterogeneity, Modality richness and imbalance, Bandwidth and compute constraints, On-device continual learning, Distributed control and autonomy, and Yielding safety, privacy, and personalization. For each, we identify concrete challenges and envision actionable research directions. We also present an evaluation framework for deploying FFMs in embodied AI systems, along with the associated trade-offs.
@@ I Introduction
Embodied AI refers to artificial intelligence systems that are physically situated in the world — typically within robots or agents that can sense, act, and learn through interaction with their environment. Building on this foundation, embodied AI is not just redefining the role of intelligent systems, it is reimagining their very nature. What fundamentally distinguishes the next generation of embodied AI agents from traditional AI systems, such as Large Language Models (LLMs) or static vision classifiers, is their demand for interactive, physically-grounded intelligence [1]. In particular, embodied agents must continuously perceive the world through multiple sensor modalities (e.g., vision, touch, audio), interact with dynamic environments, and adapt to diverse tasks, from object manipulation and social interaction to navigating unstructured terrain for search-and-rescue and assisting surgeons in hospitals.
@@
These requirements cannot be met by the deployment of narrowly trained, single-task models at these AI agents. Instead, they naturally align with the capabilities of emerging multi-modal multi-task (M3T) foundation models (FMs), which are large-scale architectures often pretrained on diverse datasets that span language, visual scenes, and human instructions [2]. M3T FMs can provide a unified semantic backbone for embodied agents, enabling them to interpret instructions, understand environments, and plan actions. For example, a kitchen robot could leverage a single M3T FM to recognize ingredients, follow verbal commands, and manipulate utensils, even in settings it has never seen (e.g., through few/zero-shot learning).
@@
Nevertheless, applying FMs to embodied AI introduces new challenges that call for a migration from their conventional centralized training/fine-tuning. In essence, each robot/agent experiences the world through its own embodiment: different sensors, actuators, tasks, and user interactions. Further, these robots/agents operate in decentralized physical environments (e.g., homes, hospitals, and factories), where through their embodiment they accumulate rich, contextual, and privacy-sensitive data (e.g., confidential industrial processes) that cannot be easily pooled/centralized at scale. Subsequently, to truly realize the potential of M3T FMs in these settings, we should move toward cross-embodiment learning, where embodied agents that are inherently data collectors can share, refine, and adapt M3T FMs through decentralized collaborations despite their heterogeneous configurations. Here, Federated Learning (FL) offers a compelling mechanism for this collaboration, enabling distributed agents to share model updates without transmitting raw data, thereby preserving privacy [3, 4].
@@
In this work, we introduce Federated Foundation Models (FFMs) for embodied AI, a natural yet underexplored solution to the challenges and motivations outlined above. Integration of FFMs in this domain creates a new paradigm that brings together the expressive generalization power of FMs with the privacy-preserving, decentralized adaptation/learning capabilities of FL. To give our discussions a unified theme, we identify the most relevant aspects of embodied AI that affect the implementation of FFMs over the network edge under EMBODY dimensions: Embodiment heterogeneity (hardware, sensors, actuators), Modality richness and imbalance, Bandwidth and compute constraints, On-device continual learning, Distributed control and autonomy, Yielding safety, privacy, and personalization. This work is created with the purpose of being a vision paper that both illuminates the transformative potential of integrating FFMs in embodied AI and expose the key challenges arising from such integration. Our contributions are summarized below.
@@
  • We propose an architecture for FFMs suitable for embodied AI, featuring modular sensor encoders, Mixture-of-Experts (MoE) layers, and task heads.
@@
  • We concretize the EMBODY dimensions, showcase the unique capabilities of FFMs for embodied AI, and highlight various use cases of FFMs in this domain.
@@
  • We outline various research directions grounded in the EMBODY dimensions, unveiling the unique opportunities that the modularity of FFMs offer for embodied AI. These directions are intentionally framed to capture the broader theme of “what can be done” in this underexplored research area, offering a flexible conceptual basis for future decomposition into specific, actionable research studies.
@@
  • We envision an evaluation framework for FFMs in embodied AI, consisting of different evaluation metrics and tradeoffs.
TABLE I: Comparative analysis of FL-only, FM-only, and FFM approaches in the embodied AI domain.
Dimension FL-only Robotics Systems FM-only Robotics Systems Federated Foundation Models (FFMs)
Modality Handling Typically uni- or bi-modal (e.g., vision + depth); often fixed during the training Trained on massive cross-modal corpora; supports rich modality fusion Supports rich, time-varying multi-modal input (vision, haptics, and audio) via modular encoders and shared latent representation
Training Setup Typically distributed training from scratch Centralized pretraining; often uses large-scale datasets Decentralized cooperative modular training; enables continual modular modifications
Personalization Local model adaptation or clustering-based personalization Typically lacks per-agent personalization; assumes centralized adaptation Lightweight personalization via on-device modules, prompt tuning, or adapters while preserving global backbone consistency
Generalization Across Embodiments Weak; performance drops across agent types or sensor configurations Moderate; limited adaptability to embodiment diversity and emerging modalities/tasks without centralized retraining/fine-tuning Strong; supports embodiment-conditioned learning, modular control heads, and cross-agent module transfers
Scalability Across Agents Designed for multi-agent participation Not designed for multiple-agent adaptation or training Designed for multi-agent participation with sparse, asynchronous updates and scalable modular coordination mechanisms
Privacy Guarantees Strong; data remains on-device; suitable for sensitive environments Weak; requires centralized data ingestion and fine-tuning Strong; Maintains privacy through federated aggregation while keeping the raw data local
Adaptation Frequency Episodic or periodic retraining Offline adaptation; large update intervals; not suitable for in-the-loop updates Real-time or task-triggered on-device adaptation via modular updates (e.g., encoders, adapters, prompt tuning)
Update Efficiency Lightweight updates due to the use of small models but often low expressiveness Efficient modular, adapter, or prefix/prompt-based updates, centralized fine-tuning required; privacy-risk Efficient modular, adapter, or prefix/prompt-based updates suitable for distributed privacy-preserving edge deployment
Safety and Interpretability Easier to interpret due to model simplicity; can be fine-tuned to obey the safety measures; lacks generalization Hard to interpret; Modular components support targeted auditing and safe rollback of local adaptations; emergent behavior may violate safety unless centrally verified Hard to interpret; Modular components support distributed on-device targeted auditing and safe rollback of local adaptations
Prompt Tuning Support Rarely supported or relevant due to small model scale Centralized prompt tuning enables efficient adaptation to new tasks and sensor conditions without modifying core weights Distributed on-device prompt tuning enables efficient adaptation to new tasks and sensor conditions without modifying core weights
@@
Figure 1: (a): A schematic of FL architecture, where clients/agents engage in collaborative model training. The collaborative training occurs through the repetition of (i) local model training at the agents using their local data, (ii) model/gradient transmission to the server, and (iii) updating the global model at the server based on the received models/gradients (e.g., via weighted averaging) and the broadcast of the global model from the server to the agents to initiate the next round of local model training. (b): Modular architecture of M3T FMs, comprising modality encoders, MoMEs, MoTEs, and task heads. Based on the input characteristics, a subset of MoMEs are triggered/engaged in the inference and training. Further, based on the desired output tasks, a subset of MoTEs will be activated. (c): Architecture of FFMs in embodied AI, where different agents have different modalities and tasks of interest. Each agent possesses a local FM, trains different modules (e.g., encoder, task head, subset of MoMEs or MoTEs) of its local FM, and transmits them to the server for aggregation. The server aggregates the received modules and broadcasts these aggregated modules back to the agents. The received aggregated modules can further go through a local fine-tuning at the agents.
@@ II Background and Related Work
II-A Embodied AI
Modern instances of embodied AI agents are humanoid robots (e.g., Figure 01 by Figure AI), mobile robotic platforms (e.g., Boston Dynamics’ Spot), and extended reality (XR) agents operating through immersive wearables, such as Meta Quest or Apple Vision Pro [1, 5]. Unlike traditional AI systems, mostly confined to screen-based tasks, embodied agents interact with dynamic, uncertain physical and social environments.
Embodied agents fundamentally depend on multi-modal perception to gain an understanding of their surroundings through a diverse range of sensory data, such as visual (RGB, depth), auditory, tactile, inertial, and haptic, often in real-time and under varying quality conditions. Also, to enable goal-directed behaviors, these agents may need to tightly couple perception with motor actions, comprising walking, turning, or navigating (i.e., locomotion), grasping or moving objects (i.e., manipulation), gesturing or signaling (i.e., expressive movement), and operating tools or interfaces. Such motor actions should often be performed under strict latency constraints to respond to environmental or user-driven changes. Nevertheless, unlocking the use of the multi-modal perceived data of these agents to train machine learning (ML) models capable of guiding motor actions under strict latency constraints is not a trivial task: these agents, such as home robots or wearable XR systems, collect deeply personal data (e.g., household layouts, gestures, voice), which cannot be transferred to centralized ML model training units (e.g., cloud servers) due to privacy concerns. This motivates the introduction of privacy-preserving distributed ML realized by FL to embodied AI.
II-B FL in Robotics/Embodied AI
@@
Conventional FL (see Fig. 1(a)) operates through a repeated three-step process until model convergence: (i) each client/device trains a local model using its data; (ii) model updates, such as parameters/gradients, are periodically shared with a server/aggregator; and (iii) the aggregator combines these updates (e.g., via weighted averaging) to a global model and broadcasts it to the devices, thereby synchronizing their local models and initiating the next round of training. FL has been applied in robotics and embodied AI for tasks such as cooperative driving and collaborative manufacturing, enabling robots to jointly learn motion plans and safety-critical controls. Its variants, such as decentralized/gossip-based FL, federated reinforcement learning (FRL), and blockchain-based FL, have also been explored. For example, gossip-based FL has been used in cobot systems, while FRL is used for multi-agent policy learning. For a detailed discussion, see [3, 4]. Collectively, these advances underline FL’s role in scalable intelligence for embodied AI; however, these works do not focus on the FMs.
II-C FMs for Embodied AI
@@
FMs have evolved rapidly, starting with single-modal LLMs (e.g., BERT, GPT-3, and PaLM), showcasing strong text generation/comprehension [6]. This was followed by multi-modal FMs, such as CLIP, DALL-E, and Imagen, enabling zero-shot classification and text-to-image generation [7]. Recently, M3T FMs (e.g., Gato, Kosmos-2, and GPT-4) have emerged, aiming for general-purpose AI that learns/reasons/acts across diverse tasks and modalities using unified architectures [8]. Although FMs are new to embodied AI, some pioneering works exist: RT-1 [9] trained a transformer on 130k robot trajectories; RoboCat [10] explored cross-embodiment learning; and SayCan [11] and ChatGPT-for-Robotics [12] applied LLMs for planning and code generation. These studies show the promise of FMs in embodied AI but assume centralized training and overlook the modularity of modern M3T FMs, an aspect we will explore in the context of FFMs.
II-D FFMs for Distributed Embodied AI
@@
FFMs, especially when considering the emerging M3T FFMs, are a highly recent research topic. Subsequently, they are quite unexplored in embodied AI. Nevertheless, recent research on FFMs in other domains has shown their tremendous potential, and the study of their aggregation methods, computational/communication efficiency, and trustworthiness is gaining substantial interest in the AI/ML and wireless networking community (refer to [13] and the references therein). In this work, we aim to provide one of the first visions for the integration of M3T FFMs in the embodied AI domain. To provide a more structured understanding, in Table I, we provide a comparative analysis of three major paradigms in embodied AI: FL-only systems, FM-only systems, and our proposed FFMs.
II-E Modular Architecture of M3T FFMs
@@
There is no unified architecture for M3T FMs as they are still under active development and envisioned differently across tech companies and academic literature. In this work, we build upon the proposal architecture in [14] and decompose it into a modular architecture, which is depicted in Fig. 1(b), where updates/training can be applied to its various modules independently. Subsequently, to enhance the comprehension of the M3T FFMs, we put them into the context of embodied AI, as illustrated in Fig. 1(c), and explain their components below.
@@
5. Coordinator and Learning Process: A central coordinator/server manages model updates across agents, following the standard FL process (i.e., local training, aggregation, and broadcast) without exchanging full-model parameters. Here, agents update only local FM sub-components/modules (e.g., encoders, heads, adapters, expert weights, prompts), whose selection can be optimized as we discuss later in the future research directions. The coordinator aggregates these modules and returns them to agents for further updates.
$\star$ Henceforth, we use ‘FFM’ to refer to ‘M3T FFM’ since our focus is solely on this type.
Figure 2: Visualizations of the envisioned future research directions. (a): Embodiment Heterogeneity. Left Plot: Embodiment-Conditioned Expert Routing. Right Plot: Cross-Embodiment Consistency Regularization. (b): Modality Richness and Imbalance. Left Plot: Modality-Aware Clustered Training. Right Plot: Cross-Modal Recovery via Expert Cascades. (c): Bandwidth and Compute Constraints. Left Plot: Context-Aware Federated Caching. Right Plot: Context-Aware and Event-Triggered Scheduling of FFM Modules.

### v2

@@ Abstract
As embodied AI systems become increasingly multi-modal, personalized, and interactive, they must learn effectively from diverse sensory inputs, adapt continually to user preferences, and operate safely under resource and privacy constraints. These challenges expose a pressing need for machine learning models capable of swift, context-aware adaptation while balancing model generalization and personalization. Here, two methods emerge as suitable candidates, each offering parts of these capabilities: multi-modal multi-task foundation models (M3T-FMs) provide a pathway toward generalization across tasks and modalities, whereas federated learning (FL) offers the infrastructure for distributed, privacy-preserving model updates and user-level model personalization. However, when used in isolation, each of these approaches falls short of meeting the complex and diverse capability requirements of real-world embodied AI environments. In this vision paper, we introduce multi-modal multi-task federated foundation models (M3T-FFMs) for embodied AI, a new paradigm that unifies the strengths of M3T-FMs with the privacy-preserving distributed training nature of FL, enabling intelligent systems at the wireless edge. We collect critical deployment dimensions of M3T-FFMs in embodied AI ecosystems under a unified framework, which we name “EMBODY”: Embodiment heterogeneity, Modality richness and imbalance, Bandwidth and compute constraints, On-device continual learning, Distributed control and autonomy, and Yielding safety, privacy, and personalization. For each, we identify concrete challenges and envision actionable research directions. We also present an evaluation framework for deploying M3T-FFMs in embodied AI systems, along with the associated trade-offs. Finally, we present a prototype implementation of M3T-FFMs and evaluate their energy and latency performance. To foster further research in this largely untapped area, we share our implementation through an open-source repository (GitHub: https://github.com/payamsiabd/M3T-FFM-EmbodiedAI).
@@ I Introduction
Embodied AI refers to artificial intelligence systems that are physically situated in the world — typically within robots or agents that can sense, act, and learn through interaction with their environment, e.g., “Figure 01” by Figure AI, “Boston Dynamics’ Spot”, and “Meta Quest” or “Apple Vision Pro” extended reality (XR) devices [1, 2]. Embodied AI is not just redefining the role of intelligent systems, it is reimagining their very nature. What fundamentally distinguishes the next generation of embodied AI agents from traditional AI systems, such as Large Language Models (LLMs) or static vision classifiers, is their demand for interactive, physically-grounded intelligence. In particular, embodied agents must continuously perceive the world through multiple sensor modalities (e.g., vision, touch, audio), interact with dynamic environments, and adapt to diverse tasks, from object manipulation and social interaction to navigating unstructured terrain for search-and-rescue and assisting surgeons in hospitals.
@@
These requirements cannot be met by the deployment of narrowly trained, single-task models at these AI agents. Instead, they naturally align with the capabilities of emerging multi-modal multi-task (M3T) foundation models (FMs), which are large-scale architectures often pretrained on diverse datasets that span language, visual scenes, and human instructions [3]. M3T-FMs can provide a unified semantic backbone for embodied agents, enabling them to interpret instructions, understand environments, and plan actions. For example, a kitchen robot could leverage a single M3T-FM to recognize ingredients, follow verbal commands, and manipulate utensils, even in settings it has never seen (e.g., through few/zero-shot learning).
@@
Nevertheless, applying M3T-FMs to embodied AI introduces new challenges that call for a migration from their conventional centralized training/fine-tuning. In essence, each robot/agent experiences the world through its own embodiment: different sensors, actuators, tasks, and user interactions. Further, these robots/agents operate in decentralized physical environments (e.g., homes, hospitals, and factories), where through their embodiment they accumulate rich, contextual, and privacy-sensitive data (e.g., confidential industrial processes) that cannot be easily pooled/centralized at scale. Subsequently, to truly realize the potential of M3T-FMs in these settings, we should move toward cross-embodiment learning, where embodied agents that are inherently data collectors can share, refine, and adapt M3T-FMs through decentralized collaborations despite their heterogeneous configurations. Here, Federated Learning (FL) offers a compelling mechanism for this collaboration, enabling distributed agents to share model updates without transmitting raw data, thereby preserving privacy [4, 5].
@@
In this work, we introduce M3T Federated Foundation Models (FFMs) for embodied AI, a natural yet underexplored solution to the challenges and motivations outlined above. Integration of M3T-FFMs in this domain creates a new paradigm that brings together the expressive generalization power of M3T-FMs with the privacy-preserving, decentralized adaptation/learning capabilities of FL. To give our discussions a unified theme, we identify the most relevant aspects of embodied AI that affect the implementation of M3T-FFMs over the network edge under EMBODY dimensions: Embodiment heterogeneity (hardware, sensors, actuators), Modality richness and imbalance, Bandwidth and compute constraints, On-device continual learning, Distributed control and autonomy, Yielding safety, privacy, and personalization. This work is created with the purpose of being a vision paper that both illuminates the transformative potential of integrating M3T-FFMs in embodied AI and expose the key challenges arising from such integration. Our contributions are summarized below.
@@
  • We propose an architecture for M3T-FFMs suitable for embodied AI, featuring modular sensor encoders, Mixture-of-Experts (MoE) layers, and task heads.
@@
  • We concretize the EMBODY dimensions, showcase the unique capabilities of M3T-FFMs for embodied AI, and highlight various use cases of M3T-FFMs in this domain.
@@
  • We outline various research directions grounded in the EMBODY dimensions, unveiling the unique opportunities that the modularity of M3T-FFMs offer for embodied AI. These directions are intentionally framed to capture the broader theme of “what can be done” in this underexplored research area, offering a flexible conceptual basis for future decomposition into specific, actionable research studies.
@@
  • We envision an evaluation framework for M3T-FFMs in embodied AI, consisting of different evaluation metrics and tradeoffs.
TABLE I: Comparative analysis of FL, M3T-FMs, and M3T-FFM approaches in the embodied AI domain.
@@
Figure 1: (a): A schematic of FL architecture, where clients/agents engage in collaborative model training. The collaborative training occurs through the repetition of (i) local model training at the agents using their local data, (ii) model/gradient transmission to the server, and (iii) updating the global model at the server based on the received models/gradients (e.g., via weighted averaging) and the broadcast of the global model from the server to the agents to initiate the next round of local model training. (b): Modular architecture of M3T-FMs, comprising modality encoders, MoMEs, MoTEs, and task heads. Based on the input characteristics, a subset of MoMEs are triggered/engaged in the inference and training. Further, based on the desired output tasks, a subset of MoTEs will be activated. (c): Architecture of FFMs in embodied AI, where different agents have different modalities and tasks of interest. Each agent possesses a local FM, trains different modules (e.g., encoder, task head, subset of MoMEs or MoTEs) of its local FM, and transmits them to the server for aggregation. The server aggregates the received modules and broadcasts these aggregated modules back to the agents. The received aggregated modules can further go through a local fine-tuning at the agents. (d): Alternative model backbone structures in FMs comprising of stacked transformers (e.g., ChatGPT), multi-encoder fusion (e.g., CLIP), VauLT, and Flamingo (with frozen pre-trained LLM blocks and gated attention blocks trained from scratch).
@@ II Background and Related Work
II-A FL in Robotics/Embodied AI
@@
Conventional FL (see Fig. 1(a)) operates through a repeated three-step process until model convergence: (i) each client/device trains a local model using its data; (ii) model updates, such as parameters/gradients, are periodically shared with a server/aggregator; and (iii) the aggregator combines these updates (e.g., via weighted averaging) to a global model and broadcasts it to the devices, thereby synchronizing their local models and initiating the next round of training. FL has been applied in robotics and embodied AI for tasks such as cooperative driving and collaborative manufacturing, enabling robots to jointly learn motion plans and safety-critical controls [4, 5]. These applications particularly benefit from FL’s few-shot learning capabilities and its ability to generalize across diverse environments and robot embodiments [6]. Collectively, these advances underline FL’s role in scalable intelligence for embodied AI; however, these works do not focus on the FMs.
II-B FMs for Embodied AI
@@
FMs have evolved rapidly, starting with single-modal LLMs (e.g., GPT-3), followed by multi-modal FMs, such as DALL-E. More recently, M3T-FMs (e.g., GPT-4) have emerged, aiming for general-purpose AI that learns/reasons/acts across diverse tasks and modalities. Although FMs are new to embodied AI, some pioneering works exist: RT-1 [7] trained a transformer on robot trajectories; RoboCat [8] explored cross-embodiment learning; SayCan [9] and ChatGPT-for-Robotics [10] applied LLMs for planning and code generation; UniAct [11] studied FM-driven actions for embodied AI agents; ECBench [12] proposed a benchmark to evaluate FM-enabled embodied AI agents. These studies show the promise of FMs in embodied AI but assume centralized training and overlook the modularity of modern M3T-FMs, aspects that we explore in this work.
II-C FFMs for Distributed Embodied AI
@@
FFMs, especially when considering the emerging M3T-FFMs, are a highly recent research topic. Subsequently, they are quite unexplored in embodied AI. Nevertheless, recent research on FFMs in other domains has shown their tremendous potential, and the study of their aggregation methods and computational/communication efficiency is gaining substantial interest [13]. In this work, we aim to provide one of the first visions for the integration of M3T-FFMs in the embodied AI domain. To provide a structured understanding, in Table I, we compare FL, M3T-FMs, and M3T-FFMs in embodied AI.
II-D M3T-FMs Modular Architecture and M3T-FFMs Operations
@@
There is no unified architecture for M3T-FMs as they are still under active development and envisioned differently across tech companies and academic literature. In this work, we build upon the proposal architecture in [14] and decompose it into a modular architecture, which is depicted in Fig. 1(b), where updates/training can be applied to its various modules independently. Subsequently, to enhance the comprehension of the M3T-FFMs, we put them into the context of embodied AI, as illustrated in Fig. 1(c), and explain their components below.
@@
We note that the model backbone can follow more conventional architectures, such as stacked transformers, multi-encoder fusion, VauLT, and Flamingo, which are depicted at the bottom of Fig. 1.
@@
5. Coordinator and Learning Process: A central coordinator/server manages model updates across agents, following the standard FL process (i.e., local training, aggregation, and broadcast) without exchanging full-model parameters. Here, agents update only local FM sub-components/modules (e.g., encoders, heads, adapters, expert weights, prompts), whose selection can be optimized as we discuss later in the future research directions. The coordinator aggregates these modules and returns them to agents for further updates. Specifically, the aggregation process mirrors that of conventional FL, with the key distinction that it operates at the module level rather than the entire model (e.g., through weighted averaging of module parameters or alternative forms of inter-module knowledge sharing [14]). This modular update scheme enables flexible module coordination strategies tailored to each module’s role and dynamics. For instance, subsets of expert/MoE modules may undergo asynchronous, low-frequency aggregation if they already exhibit a high performance. Conversely, task heads associated with emerging or rapidly evolving tasks may benefit from frequent, synchronous updates for accelerated convergence.
$\star$ Henceforth, we use ‘FFM’ to refer to ‘M3T-FFM’ since our focus is solely on this type.
Figure 2: Visualizations of the envisioned future research directions. (a): Embodiment Heterogeneity. Left Plot: Embodiment-Conditioned Information Routing. Right Plot: Cross-Embodiment Consistency Regularization. (b): Modality Richness and Imbalance. Left Plot: Modality-Aware Clustered Training. Right Plot: Cross-Modal Recovery via Module Cascades. (c): Bandwidth and Compute Constraints. Left Plot: Context-Aware Module Caching. Right Plot: Event-Triggered Module Scheduling.